OpenStudy (anonymous):

Use the Law of the Lever to solve the following. Mark sits 4.5 feet from the center of the seesaw to balance a friend who weighs 53 lbs and is sitting 6 feet from the center. How many pounds does Mark weigh? Round to the nearest hundredth. A. 39.75 lb B. 70.67 lb C. 55.78 lb

OpenStudy (anonymous):

You should get 53*6=4.5*m,where m is the mass of Mark's friend. So, m= 53*6/4.5 = 70.67 lb to the nearest hundredth.